OpenStudy (anonymous):

When rewriting (x^2 +2x + x^1/3) / x^2/3 to have the denominator moved to the numerator. Will the x^-2/3 be multiplied to the numerator or added?

OpenStudy (aum):

Multiplied. \[\large \frac{(x^2 +2x + x^{1/3})}{ x^{2/3}} = x^{-2/3} \times (x^2 +2x + x^{1/3}) \]Now you can distribute the \(\large x^{-2/3}\). When multiplying, since the bases are the same, the exponents should be added.